# MaskMoE: Boosting Token-Level Learning via Routing Mask in Mixture-of-Experts

### v1

@@ Abstract
Scaling model capacity enhances its capabilities but significantly increases computation. Mixture-of-Experts models (MoEs) address this by allowing model capacity to scale without substantially increasing training or inference costs. Despite their promising results, MoE models encounter several challenges. Primarily, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens. Additionally, while fixed routing mechanisms can mitigate this issue, they compromise on the diversity of representations. In this paper, we propose MaskMoE, a method designed to enhance token-level learning by employing a routing masking technique within the Mixture-of-Experts model. MaskMoE is capable of maintaining representation diversity while achieving more comprehensive training. Experimental results demonstrate that our method outperforms previous dominant Mixture-of-Experts models in both perplexity (PPL) and downstream tasks.
@@ 1 Introduction
Large language models have achieved promising performance on various downstream natural language tasks Touvron et al. (2023a); Dai et al. (2022); Brown et al. (2020); Anil et al. (2023). However, training large language models on extensive textual data has significantly increased computational costs compared to previous works (e.g., BERT Devlin et al. (2019), ELMO Peters et al. (2018), and LSTM Hochreiter and Schmidhuber (1997)). In order to further scale up models within computational budgets, sparse activation networks Child et al. (2019); Du et al. (2022a) receive widespread attention due to their ability to significantly reduce computational costs by using only a few parameters per input. A widely studied approach is Mixture-of-Experts (MoEs)  Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022); Roller et al. (2021), which trains multiple expert networks but selects only a subset to process specific inputs. Compared to dense networks of the same model size, MoEs effectively reduce computational costs.
@@
The current commonly used dynamic routing method in MoEs is to select the expert with the high confidence based on the probability distribution output by an intermediate layer with learnable parameters, i.e., a softmax layer. Previous works Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022) show that MoE models trained based on dynamic routing achieve better performance than dense models with the same amount of training computation.
@@
However, this dynamic routing method introduces a new problem: the routing fluctuation problem Dai et al. (2022). This implies that as training progresses, the same tokens are assigned to different experts in different iterations. Consequently, each expert is trained on only a subset of the same tokens. Additionally, the lack of knowledge sharing and interaction among experts may negatively impact the model’s performance Zhao et al. (2024). Specifically, for the same tokens assigned to different experts, the knowledge each expert learns is harder to share. As previous work Shazeer et al. (2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher perplexity than scenarios with fewer experts. The MoEs models have relatively less impact on the learning of frequent tokens, as these tokens have sufficient data to ensure that each expert receives adequate training. However, the MoEs model disperses infrequent tokens and concepts across various experts, which can lead to underfitting for these tokens. This issue hinders scaling the number of experts for the MoE model. As the number of experts increases while keeping the training data constant, the average number of tokens assigned to each expert further decreases, exacerbating the problem of expert underfitting.
@@
Figure 1: Explanation of random routing masking strategy. For each token, some experts will be randomly masked, with only a subset of experts being visible, e.g., for the token “in” only experts 1, 2 and 3 are visible.
@@
Previously fixed routing strategies Roller et al. (2021) based on Hash Layer, pre-assigning each token to a fixed-route expert, e.g., “deliberate” can only be routed to expert number one as shown in Figure 1. The fixed routing to send the same token to the same expert is more advantageous for the token’s thorough learning, especially for some infrequent tokens. However, previous works He et al. (2023); Yang et al. (2019) show that increasing the number of optional experts for the same token input enhances the diversity of representations. Therefore, fixed routing based on the Hash Layer may represent a deficiency in diversity, especially for the representation of frequent tokens. Frequent tokens, in order to achieve diverse representations, require encoding by more experts due to their broader range of usage Koranda et al. (2018). For instance, “the” can follow a variety of nouns, and having a more diverse representation helps to distinguish these nouns, making it easier to predict them in the next token prediction task. To enhance the diversity of representation for frequent tokens, it is necessary to employ more experts to encode them.
@@
In this paper, we propose a routing strategy called MaskMoE. Specifically, we propose a routing masking method to adjust the number of visible experts for tokens of different frequencies by generating a masking vector for each token in the vocabulary before training. For infrequent tokens, the MoE layers employ routing masking to retain only one visible expert to which the token can be routed. For frequent tokens, the MoE layers have a lower routing masking rate, allowing more visible experts. MaskMoE can enhance token-level learning through a routing mask. Specifically, MaskMoE allows the model to learn more intensively about infrequent tokens by routing the same infrequent tokens to the same expert, enabling the expert to be fully trained on those tokens. Meanwhile, MaskMoE maintains the diversity of representations for frequent tokens by still having multiple experts available for routing, which benefits their representation diversity He et al. (2023); Yang et al. (2019). Additionally, despite frequent tokens being routed to multiple experts, their sufficient quantities ensure thorough training. Increase the visible number of experts for high-frequency tokens, i.e., increasing the count of updatable parameters, can further enhance capability. Our experimental results indicate that MaskMoE outperforms previous MoEs models, whether in terms of perplexity or downstream benchmark datasets.
@@
  • We propose MaskMoE, which introduces a routing mask method to assign a different number of visible experts to tokens based on their frequency to enhance the model’s token-level learning. MaskMoE ensures sufficient training for infrequent tokens while maintaining diverse representations for frequent tokens.
  • We conduct extensive experiments to analyze the relationship between the number of experts and token frequency. We highlight that dynamic routing, which disperses the same tokens, can lead to underfitting of experts on infrequent tokens, while fixed routing lacks representation diversity for frequent tokens.
@@ 2.1 Language Models
Language models are statistical models designed to optimize the likelihood of token sequences in training data Touvron et al. (2023a). Initially, language models relied on $n$-gram statistics Bahl et al. (1983); Katz (1987); Kneser and Ney (1995). Subsequently, the emphasis shifted to neural network-based models, particularly Recurrent Neural Networks Mikolov et al. (2010) and their variants like LSTMs Graves (2013). These models have demonstrated the ability to learn intricate patterns within textual data, achieving significant success in various language modeling applications.
In recent times, Transformers have become the predominant architecture for language models. Notable examples include BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), UniLM Dong et al. (2019), and T5 Raffel et al. (2020). The introduction of GPT-3 Brown et al. (2020), which boasts 175 billion parameters, marked a significant milestone due to its exceptional performance across numerous downstream tasks. This has led to a surge in research focusing on large generative language models, including prominent works such as Gopher Rae et al. (2021), PaLM Chowdhery et al. (2022), Pythia Biderman et al. (2023), OPT Zhang et al. (2022), GLM Du et al. (2022b); Zeng et al. (2023) and LLaMA Touvron et al. (2023a, b). Currently, GPT4 OpenAI (2023), trained with Reinforcement Learning from Human Feedback (RLHF), achieves truly remarkable results.
However, as the size of the model grows, the computational demands for both training and inference also increase. MoE models achieve scalability by sparsely activating portions of the model’s parameters, allowing for an increase in model size without significantly raising computational overhead. Consequently, MoE models have been receiving increasing attention recently.
@@ 2.2 Mixture-of-Experts
The concept of Mixture-of-Experts (MoE) models is initially proposed by  Jacobs et al. (1991). Later,  Shazeer et al. (2017) applies MoE to LSTM, training an LSTM model with up to 137B parameters. With the rise of the Transformer architecture Vaswani et al. (2017); Devlin et al. (2019), Gshard Lepikhin et al. (2021) first applies MoE to Transformers. Subsequently, powerful MoE models such as GLaM Du et al. (2022a) and Switch Transformer Fedus et al. (2022) emerge.
Early works  Zoph et al. (2022); Fedus et al. (2022); Du et al. (2022a); Lepikhin et al. (2021) mainly focus on dynamic routing methods for learning-to-route MoEs. The Hash Layer  Roller et al. (2021) proposes using a random hashing method to route tokens to a fixed expert, achieving better results than dynamic routing. Then,  Dai et al. 2022 points out that dynamic routing has the problem of routing fluctuation problem, meaning the same input is assigned to different experts as training progresses. Routing fluctuation often harms sample efficiency because the same input updates different experts, which significantly affects infrequent tokens. However, previous works He et al. (2023); Yang et al. (2019) indicate that having more experts allows tokens to obtain richer representations, and fixed routing can affect the diversity of representations, especially for frequent tokens.
Based on the above analysis, we propose MaskMoE, which uses routing masks to alleviate the underfitting problem of infrequent tokens caused by routing fluctuations, while maintaining the representational diversity of frequent tokens.
@@ 3 Method
In this section, we first review the training of the MoEs models. Then, we introduce detailed MaskMoE. Finally, we introduce the load balance loss.
@@ 3.1 Reviewing the Mixture-of-Experts
The routing of MoEs can be divided into fixed routing Roller et al. (2021); Dai et al. (2022) and dynamic routing with learnable parameters Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022). The common method for fixed routing is to use random hashing to determine Roller et al. (2021), before training, which experts each token will be sent to. In contrast, dynamic routing employs a routing layer with learnable parameters, i.e., a softmax layer, to decide which expert should process the input tokens.
In terms of structure, the FFN layer of the MoEs model is replaced with sparse expert modules. Given a tokenized input sequence $\mathbf{x}={x_{1},x_{2},...,x_{T}}$ consisting of $T$ tokens, the representation of each token is computed as hidden states by a standard Transformer.
$\mathbf{H}=Tansformer(Embedding(\mathbf{x}))$ (1)
Here, $\mathbf{H}$ denote the hidden states, i.e., $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\ldots,\mathbf{h}_{T}]$. The dense transformer layer consists of self-attention and FFN sublayers.
$\mathbf{\bar{h}}_{t}^{l}=SelfAttn(\mathbf{h_{t}^{l-1}})$ (2)
$\mathbf{h_{t}^{l}}=FFN(\mathbf{\bar{h}}_{t}^{l})$ (3)
Here, $l$ represents the $l$-th transformer layer. MoEs methods replace the FFN sub-layer with the MoEs module. Each MoEs module consists of multiple FFNs, whose outputs are mixed with the gating function $P(\cdot)$, represented by the following formula.
$\mathbf{h_{t}^{l}}=\sum_{i}^{N}P(\mathbf{\bar{h}}_{t}^{l})\cdot FFN_{i}(\mathbf{\bar{h}}_{t}^{l})$ (4)
$N$ represents the number of experts in a single MoE layer. The majority of elements in the distribution of the gated $P(\cdot)$ output are zeros, i.e., only a small portion of experts will be activated. Therefore, increasing the total number of experts in MoEs does not increase computation or inference time.
@@ 3.2 Proposed MaskMoE
The typical dynamic routing method routes tokens to the top experts based on the scores output by a softmax layer. This approach to update expert parameters is sparse, meaning that the parameters of unselected experts are not updated. This sparsity may lead to experts underfitting for infrequent tokens. Fixed routing ensures that specific experts consistently process the same input tokens, which helps mitigate the problem of insufficient learning for infrequent tokens. However, it is less effective in fostering a diversity of representations compared to dynamic routing methods. Dynamic routing allows the same token to be encoded by different experts, leading to more varied representations He et al. (2023); Yang et al. (2019). Since frequent tokens can have different meanings depending on the context, fixed routing is less effective than dynamic routing in capturing the nuances of these tokens. To balance the density of parameter updates and the diversity of representations in MoE models, we propose MaskMoE to boost token-level learning via routing mask. As illustrated in Figure 1, MaskMoE achieves this by introducing a routing masking mechanism. For each token, a fixed random masking is applied to the experts, allowing them to be routed only to their respective subsets of experts. The formula is represented as follows.
$\mathbf{p}$ $=\text{Softmax}(W_{g}\mathbf{\bar{h}_{t}^{l}}+\mathbf{M})$ (5)
Where $\mathbf{M}$ is the masking vector used to control the visibility of token-specific experts. For visible experts, the corresponding elements in $\mathbf{M}$ are 0, while for invisible experts, the corresponding elements in $\mathbf{M}$ are $\infty$- ∞. The masking vector is determined before training and does not change during the training process. For multi-layer MoE models, we reuse the same masking vector $\mathbf{M}$ across different MoE layers. Here, $\mathbf{W_{g}}$ is the parameter of the gate. The initial energy masking vector for a token $t$ can be represented by the following formula.
@@
Formally, we denote the number of visible experts as $V$. $\mathcal{U}(\{0:N-1\})$ denotes a uniform distribution over the set of integers from 00 to $N-1$. It is worth noting that for infrequent tokens, each token has only one visible expert, i.e., $V=1$. This promotes more thorough learning by the selected expert for such infrequent tokens. For frequent tokens, we also reduce the number of visible experts, but the number of visible experts is greater than 1, i.e., $1<V<=N$. For frequent tokens, it is common for $V$ to be less than $N$. We believe that an appropriately sized $V$ not only allows frequent words to have higher representation diversity compared to fixed routing but also enables more thorough training for the same token compared to dynamic routing.

### v2

@@ Abstract
Scaling the size of a model enhances its capabilities but significantly increases computation complexity. Mixture-of-Experts models (MoE) address the issue by allowing model size to scale up without substantially increasing training or inference costs. Despite their promising results, MoE models encounter several challenges. Primarily, for dynamic routing methods, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens. Additionally, while fixed routing methods can mitigate that issue, they compromise on the diversity of representations. In this paper, we propose MaskMoE, a method designed to enhance token-level learning by employing a routing masking technique within the Mixture-of-Experts model. MaskMoE is capable of maintaining representation diversity while achieving more comprehensive training. Experimental results demonstrate that our method outperforms previous dominant Mixture-of-Experts models in terms of both perplexity (PPL) and downstream task performance.
@@ 1 Introduction
Large language models have achieved promising performance on various downstream natural language tasks Touvron et al. (2023a); Dai et al. (2022); Brown et al. (2020); Anil et al. (2023). Moreover, according to the scaling law  Kaplan et al. (2020); Hoffmann et al. (2022), as the model size increases, the model’s size will continue to grow. However, for dense language models, the computational costs of continuing to scale up are substantial. In order to further scale up models within computational budgets, sparse activation networks Child et al. (2019); Du et al. (2022a) receive widespread attention due to their ability to significantly reduce computational costs by using only a part of parameters per input. A widely studied approach is Mixture-of-Experts (MoE)  Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022); Roller et al. (2021), which trains multiple expert networks but selects only a subset to process specific inputs. Compared to dense networks of the same model size, MoE effectively reduces computational costs and achieves better results in both PPL and downstream tasks Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024).
@@
In MoE models, for each input (e.g., a token in the language models), the router in front of the experts needs to decide which tokens to distribute to which experts. The current commonly used dynamic routing method in MoE is to select the expert with the top $k$ highest confidence based on a probability distribution output by an intermediate layer with learnable parameters. Previous works Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022) show that MoE models trained based on dynamic routing achieve better performance than dense models with the same amount of training computation.
@@
However, this dynamic routing method introduces a new challenge: the routing fluctuation problem Dai et al. (2022). It implies that as training moves on, the same token is assigned to different experts in different iterations. This has two negative impacts. Firstly, each expert is trained on only a subset of the occurrences of an identical token, resulting in each expert not learning enough about this token. As previous work Shazeer et al. (2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher PPL than its counterparts with fewer experts. The routing fluctuation problem has relatively less impact on the learning of frequent tokens, as those tokens have sufficient occurrences to ensure that each expert receives adequate training. On the contrary, for infrequent tokens, the routing fluctuation problem can disperse each of them across various experts, which can lead to underfitting for them. Secondly, the knowledge each expert learns from the corresponding occurrence subset w.r.t an identical token is hard to share among experts. Such lack of knowledge sharing and interaction among experts may negatively impact the model’s performance Zhao et al. (2024). The above two issues hinder scaling the number of experts for an MoE model. Because, as the number of experts increases while keeping the training data constant, the average number of tokens assigned to each expert further decreases, exacerbating the above two issues.
@@
Figure 1: Illustration of random routing masking method. For each token, some experts will be randomly masked, with only a subset of experts being visible, e.g., for the token “in” only expert FFN #1, #2, and #3 are visible.
@@
To alleviate the issue of routing fluctuations, the previously proposed fixed routing method Roller et al. (2021); Dai et al. (2022) is a solution. Specifically, Hash Layer Roller et al. (2021) pre-assigns the same token to a fixed routing expert. For instance, as shown in Figure 1, the token “deliberate” can only be routed to the expert FFN#2. The fixed routing method to send all the occurrences of a token to the same expert is advantageous for the token’s thorough learning, especially for infrequent tokens. However, fixed routing based on the Hash Layer represents a deficiency in diversity, especially for those frequent tokens. Previous works He et al. (2023); Yang et al. (2019) show that increasing the number of optional experts for all occurrences of an identical can enhance the diversity of representations. Frequent tokens probably require to be encoded by more experts due to their broader range of usage Koranda et al. (2018). For instance, the token “the” appears in various contexts and owns different semantic meanings, and having a more diverse representation helps to distinguish it. Therefore, based on the observations above, we need a fixed routing strategy for infrequent tokens, while also needing more experts for frequent ones.
@@
To meet the requirements above. In this paper, we propose a routing method called MaskMoE, which includes a routing masking method that adjusts the number of visible experts for tokens of different frequencies, by generating a masking vector for each token in the vocabulary before training. Specifically, for each infrequent token, the proposed MaskMoE employs routing masking to retain only one visible expert to which the token can be routed, in each MoE layer. For each frequent token, the proposed MaskMoE has a lower routing masking rate to allow more visible experts for it in each MoE layer. MaskMoE can enhance token-level learning through such a routing mask. Specifically, MaskMoE makes the model learn more intensively about infrequent tokens by routing each to an identical expert in each MoE layer, enabling the expert to be fully trained w.r.t that token. Meanwhile, MaskMoE maintains the diversity of representations for frequent tokens by still having multiple experts available for routing, which benefits their representation diversity He et al. (2023); Yang et al. (2019). Despite frequent tokens being routed to multiple experts, their sufficient quantities can still ensure thorough training. Our experimental results indicate that MaskMoE outperforms previous MoE models, in terms of either PPL or downstream task performance.
@@
  • We propose MaskMoE, which introduces a routing masking method to assign different numbers of visible experts to tokens based on their frequency to enhance token-level learning. MaskMoE ensures sufficient training for infrequent tokens while maintaining diverse representations for frequent tokens.
  • We highlight that dynamic routing, which disperses the occurrences of a token into different experts during training, can lead to the underfitting of experts on infrequent tokens, while fixed routing lacks representation diversity for frequent tokens.
@@ 2.1 Language Models
Language models are statistical models designed to optimize the likelihood of token sequences in training data Touvron et al. (2023a). Initially, language models relied on $n$-gram statistics Bahl et al. (1983); Katz (1987); Kneser and Ney (1995). Subsequently, the emphasis shifted to neural network-based models, particularly Recurrent Neural Networks Mikolov et al. (2010) and their variants like LSTMs Graves (2013). Those models have demonstrated the ability to learn intricate patterns within textual data, achieving significant success in various language modeling applications.
In recent times, Transformers have become the predominant architecture for language models. Notable examples include BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), UniLM Dong et al. (2019), and T5 Raffel et al. (2020). The introduction of GPT-3 Brown et al. (2020), which boasts 175 billion parameters, marked a significant milestone due to its exceptional performance across numerous downstream tasks. And that then led to a surge in research focusing on large generative language models, including prominent works such as Gopher Rae et al. (2021), PaLM Chowdhery et al. (2022), Pythia Biderman et al. (2023), OPT Zhang et al. (2022), GLM Du et al. (2022b); Zeng et al. (2023) and LLaMA Touvron et al. (2023a, b). Currently, GPT4 OpenAI (2023), trained with Reinforcement Learning from Human Feedback (RLHF), achieves truly remarkable results.
However, as the size of the model grows, the computational demands for both training and inference also increase. MoE models achieve scalability by sparsely activating a portion of the model’s parameters, allowing for an increase in model size without significantly raising computational overhead. Consequently, MoE models have been receiving increasing attention recently.
@@ 2.2 Mixture-of-Experts
The concept of Mixture-of-Experts (MoE) models was initially proposed by  Jacobs et al. (1991). Later,  Shazeer et al. (2017) applied MoE to LSTM, training an LSTM model with up to 137B parameters. With the rise of the Transformer architecture Vaswani et al. (2017); Devlin et al. (2019), Gshard Lepikhin et al. (2021) further applied MoE to Transformers. Subsequently, powerful MoE models such as GLaM Du et al. (2022a) and Switch Transformer Fedus et al. (2022) emerged.
Early works  Zoph et al. (2022); Fedus et al. (2022); Du et al. (2022a); Lepikhin et al. (2021) mainly focused on dynamic routing methods for learning-to-route MoE. Then the Hash Layer  Roller et al. (2021) is proposed to use a random hashing method to route tokens to a fixed expert, which even achieved better results than dynamic routing. As an explanation to the advantage of the simple fixed routing method,  Dai et al. 2022 pointed out that dynamic routing has the problem of routing fluctuation problem, meaning that an identical input is assigned to different experts as training progresses. Routing fluctuation often harms sample efficiency, especially for the learning of infrequent tokens in the language model. On the contrary, other previous works He et al. (2023); Yang et al. (2019) also argued that having more experts allows tokens to obtain richer representations, and fixed routing can affect the diversity of representations, especially for frequent tokens.
Considering both routing fluctuation and representation diversity, we propose MaskMoE, which uses routing masks to alleviate the underfitting problem of infrequent tokens caused by routing fluctuations, while maintaining the representational diversity of frequent tokens.
@@ 3 Method
In this section, we first review the training of the MoE models. Then, we elaborate on the proposed MaskMoE. Finally, we introduce the load balance loss.
@@ 3.1 Reviewing the Mixture-of-Experts
The routing of MoE can be divided into fixed routing Roller et al. (2021); Dai et al. (2022) and dynamic routing with learnable parameters Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022). The common method for fixed routing is to use random hashing to determine Roller et al. (2021), before training, which experts each token will be sent to. In contrast, dynamic routing employs a routing layer with learnable parameters to decide which expert should process an input token.
In terms of structure, the FFN layer of a dense language model is generally replaced with sparse expert modules to build an MoE model. Given a tokenized input sequence $\mathbf{x}={x_{1},x_{2},...,x_{T}}$ consisting of $T$ tokens , and then generates a probability distribution $\mathbf{p}_{v}$ over the vocabulary as output. In nearly all implementations, the Cross-Entropy loss is used as the loss function to maximize the predicted probability $\mathbf{p}_{v}^{x_{i}}$ w.r.t the ground-truth token $x_{i}$. The training loss $\mathcal{L}_{lm}$ of the generative language model can be formulated as follows.
$\mathcal{L}_{lm}$ $=-\sum_{i=2}^{T}\log(\mathbf{p}_{v}^{x_{i}})$ (1)
$s.t.,\mathbf{p}_{v}$ $=\text{softmax}(W\mathbf{H}^{L})$ (2)
$\mathbf{H^{L}}$ $=\text{Transformer}(x_{1}^{0},x_{2}^{0},\ldots,x_{T}^{0})$ (3)
Here, $\mathbf{H^{L}}=[\mathbf{h}_{1}^{L},\mathbf{h}_{2}^{L},\ldots,\mathbf{h}_{T}^{L}]$ denotes the hidden states of the last layer. With $\mathbf{H}^{L}$, a linear projection layer $W$ is introduced to derive the predicted probability distribution $\mathbf{p}_{v}$ over the vocabulary, where $\mathbf{p}_{v}=[\mathbf{p}^{x_{1}}_{v},\mathbf{p}^{x_{2}}_{v},\ldots,\mathbf{p}^{x_{T}}_{v}]$. The Transformer model consists of $L$ Transformer layers. Each layer is composed of a self-attention sublayer and an FFN sublayer, where FFN is a two-layer fully connected network.
$\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l}=SelfAttn(\mathbf{h}_{1}^{l-1},\mathbf{h}_{2}^{l-1},\ldots,\mathbf{h}_{T}^{l-1})$ (4)
$\mathbf{h}_{1}^{l},\mathbf{h}_{2}^{l},\ldots,\mathbf{h}_{T}^{l}=FFN(\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l})$ (5)
$l$ represents the $l$-th transformer layer. MoE methods replace the FFN sub-layer with the MoE module. A MoE model can replace the FFN with a MoE module for either a single layer, several layers, or even all layers. Each MoE module consists of multiple FFNs, whose outputs are mixed with the gating function $P(\cdot)$, represented by the following formula.
$\mathbf{h}_{t}^{l}=\sum_{i}^{N}P_{i}(\mathbf{\hat{h}}_{t}^{l})\cdot FFN_{i}(\mathbf{\hat{h}}_{t}^{l})$ (6)
$s.t.|P(\mathbf{\hat{h}}_{t}^{l})|_{0}<<N$ (7)
Where $N$ represents the number of experts in a single MoE layer. The majority of elements in the distribution of a gated $P(\cdot)$ output are zeros, i.e., only a small portion of experts will be activated. Therefore, increasing the total number of experts in MoE does not increase computation or inference time.
@@ 3.2 Proposed MaskMoE
Dynamic routing uses a learnable method where tokens are assigned to different experts during training, leading to routing fluctuations. On the other hand, fixed routing, like hash layers, assigns each token to a specific expert, which can also limit representation diversity. To mitigate the routing fluctuations and the diversity of representations in MoE models, we propose MaskMoE to boost token-level learning via routing mask, as illustrated in Figure 1. For each token, a fixed random masking is applied to the experts, allowing them to be routed only to their respective subsets of experts. The formula is represented as follows.
$\mathbf{p}$ $=\text{Softmax}(W_{g}\mathbf{\hat{h}_{t}^{l}}+\mathbf{m})$ (8)
Where $\mathbf{m}$ is the masking vector used to control the visibility of token-specific experts. For visible experts, the corresponding elements in $\mathbf{m}$ are 0, while for invisible experts, the corresponding elements in $\mathbf{m}$ are $\infty$- ∞. The masking vector is determined before training and does not change during the training process. For multi-layer MoE models, we reuse the same masking vector $\mathbf{m}$ across different MoE layers. Here, $W_{g}$ is the parameter of the gating function. The initial energy masking vector for a token $t$ can be represented by the following formula.
@@
We denote the number of visible experts as $V$. $N$ represents the number of experts in a single MoE layer. $\mathcal{U}(\{0:N-1\})$ denotes a uniform distribution over the set of integers from 00 to $N-1$. $\text{C}_{i}$ denotes the index of the selected expert. It is worth noting that for infrequent tokens, each token has only one visible expert, i.e., $V=1$. This promotes more thorough learning by the selected expert for such infrequent tokens. For frequent tokens, the number of visible experts is greater than 1, i.e., $1<V<=N$, and note that it is common for $V$ to be less than $N$. We believe that an appropriately sized $V$ not only allows frequent tokens to have higher representation diversity compared to fixed routing but also enables more thorough training for an identical token compared to dynamic routing. The complete process of MaskMoE is illustrated in Algorithm 1.
Algorithm 1 Proposed MaskMoE Algorithm
0:  Training dataset after tokenization $D$, the vocabulary of the tokenizer $T_{v}$, maximal training step number $S$, the number of visible experts $V$, the number of experts $N$
1:  Initialize the routing mask matrix, $\mathbf{M}\leftarrow-\infty$
2:  for each $t$ in $T_{v}$ do
3:     $\mathbf{C}=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{0:N-1\})$
4:     $\mathbf{m}^{t}=\mathbf{M}(t)$ The corresponding column w.r.t $t$
5:     $\mathbf{m}_{j}^{t}=0,\quad\forall j\in\mathbf{C}$
6:  end for
7:  for $i=1$ to $S$ do
8:     Load an input sequence $\mathbf{x}$ from $D$
9:     # Load the routing mask for each token in $\mathbf{x}$
10:     $\mathbf{m}=\mathbf{M}(\mathbf{x})$
11:     Train the MoE models with $\mathbf{x}$ as the input, and add $\mathbf{m}$ to the router according to Equation 10
12:  end for